# Diffusion approximation of the stationary distribution of a two-level single server queue

### v7

@@ 4.2 Asymptotic BAR for the pre-limit process $X^{(r)}(\cdot)$
$\varphi^{(r)}_{1}(\theta_{1})=\mathbb{E}[e^{\theta_{1}rL^{(r)}}1(L^{(r)}\leq\ell^{(r)}_{0})],\qquad\varphi^{(r)}_{2}(\theta_{2})=\mathbb{E}[e^{\theta_{2}rL^{(r)}}1(\ell^{(r)}_{0}<L^{(r)})],$
$\varphi^{(r)}_{1|cd}(\theta_{1})=\mathbb{E}[e^{\theta_{1}rL^{(r)}}|L^{(r)}\leq\ell^{(r)}_{0}],\qquad\varphi^{(r)}_{2|cd}(\theta_{2})=\mathbb{E}[e^{\theta_{2}rL^{(r)}}|\ell^{(r)}_{0}<L^{(r)}].$
@@
$w^{(r)}_{1}(z)=1(z\leq\ell^{(r)}_{0}),\qquad w^{(r)}_{2}(z)=1(z>\ell^{(r)}_{0}),\qquad z\in\mathbb{Z}_{+},$
@@
By this lemma, $g^{(r)}_{i,r\theta_{i}}(R^{(r)})$ is bounded by $e^{|\theta_{i}|(d_{e,i}+d_{s})}$ for sufficiently small $r$ for each $\theta_{1}\in\mathbb{R}$ and $\theta_{2}\leq 0$, and therefore, we have the following facts.
@@
because $r\theta_{1}L^{(r)}1(L^{(r)}\leq\ell^{(r)}_{0})\leq\theta_{1}(\ell+o(r))$ for $\theta_{1}>0$. These finiteness guarantees our computations for $\theta_{1}\in\mathbb{R}$ and $\theta_{2}\leq 0$. We also note the following fact.
@@
$\lim_{r\downarrow 0}\left|\psi^{(r)}_{i}(\theta_{i})-\varphi^{(r)}_{i}(\theta_{i}))\right|=0,\quad i=1,2.$ (4.17)
@@
$\left|\psi^{(r)}_{i}(\theta_{i})-\varphi^{(r)}_{i}(\theta_{i}))\right|$ $=\left|\mathbb{E}[e^{r\theta_{i}L^{(r)}}(g_{1,\theta_{i}}(R^{(r)}_{e})-1)1(L^{(r)}\leq\ell^{(r)}_{0})]\right|$
$\leq e^{|\theta_{1}|r\ell^{(r)}_{0}}\left|\mathbb{E}[g_{i,\theta_{i}}(R^{(r)}_{e})-1]\right|\to 0,\qquad r\downarrow 0.$
@@
Since $g^{(r)}_{i,\theta_{i}}$ changes only when $L^{(r)}(0-)=\ell^{(r)}_{0}$ under $\mathbb{P}^{(r)}_{e}$ or $L^{(r)}(0)=\ell^{(r)}_{0}$ under $\mathbb{P}^{(r)}_{s}$, (4.8) implies that
$\mathbb{E}^{(r)}_{e}\left[\Delta_{e}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0))1(L^{(r)}(0-)\not=\ell^{(r)}_{0})\right]=0,$
$\mathbb{E}^{(r)}_{s}\left[\Delta_{s}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0))1(L^{(r)}(0)\not=\ell^{(r)}_{0})\right]=0.$
Hence, from the definitions of $X^{(r)}_{1}(0)$ and $\Delta_{e}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0))$, we have
$\mathbb{E}^{(r)}_{e}[\Delta_{e}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0))]=\mathbb{E}^{(r)}_{e}[(f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}_{1}(0))-f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0-)))1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\quad=\mathbb{E}^{(r)}_{e}\Big{[}e^{r\theta_{2}(\ell^{(r)}_{0}+1)}g^{(r)}_{2,r\theta_{2}}(R^{(r)}_{1}(0))1(L^{(r)}_{1}(0)=\ell^{(r)}_{0}+1)$
$\hskip 64.58313pt-e^{r\theta_{1}\ell^{(r)}_{0}}g^{(r)}_{1,r\theta_{1}}(R^{(r)}(0-)))1(L^{(r)}(0-)=\ell^{(r)}_{0})\Big{]}$
$\quad=e^{r\theta_{2}\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s-}}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\hskip 64.58313pt-e^{r\theta_{1}\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta_{1})\widehat{R}^{(r)}_{s-}}1(L^{(r)}(0-)=\ell^{(r)}_{0}))],$
$\quad=(e^{r\theta_{2}\ell^{(r)}_{0}}-e^{r\theta_{1}\ell^{(r)}_{0}})\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s-}}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\qquad+e^{r\theta_{1}\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[(e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s-}}-e^{-\zeta^{(r)}(r\theta_{1})\widehat{R}^{(r)}_{s-}})1(L^{(r)}(0-)=\ell^{(r)}_{0}))],$ (4.22)
@@
$\mathbb{E}^{(r)}_{s}[\Delta_{s}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0))]=\mathbb{E}^{(r)}_{s}[(f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0))-f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}_{1}(0)))1(L^{(r)}(0)=\ell^{(r)}_{0})]$
$\quad=\mathbb{E}^{(r)}_{s}\Big{[}e^{r\theta_{1}\ell^{(r)}_{0}}g^{(r)}_{1,r\theta_{1}}(R^{(r)}(0))1(L^{(r)}(0)=\ell^{(r)}_{0})-e^{r\theta_{2}(\ell^{(r)}_{0}+1)}g^{(r)}_{2,r\theta_{2}}(R^{(r)}_{1}(0)))1(L^{(r)}_{1}(0-)=\ell^{(r)}_{0}+1)\Big{]}$
$\quad=e^{r\theta_{1}(\ell^{(r)}_{0}+1)}\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{1}(r\theta_{1})\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{0})]-e^{r\theta_{2}(\ell^{(r)}_{0}+1)}\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{2}(r\theta_{2})\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{0})]$
$\quad=(e^{r\theta_{1}(\ell^{(r)}_{0}+1)}-e^{r\theta_{2}(\ell^{(r)}_{0}+1)})\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{1}(r\theta_{1})\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{0})]$
$\qquad+e^{r\theta_{2}(\ell^{(r)}_{0}+1)}\mathbb{E}^{(r)}_{s}[(e^{-\eta^{(r)}_{1}(r\theta_{1})\widehat{R}^{(r)}_{e}}-e^{-\eta^{(r)}_{2}(r\theta_{2})\widehat{R}^{(r)}_{e}})1(L^{(r)}(0)=\ell^{(r)}_{0})].$ (4.23)
Define $D^{(r)}(\theta_{1},\theta_{2})=\alpha^{(r)}_{e}\left(\mathbb{E}^{(r)}_{e}\left[\Delta_{e}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0))\right]+\mathbb{E}^{(r)}_{s}\left[\Delta_{s}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0))\right]\right)$. Then, by (4.2) and (4.2),
$D^{(r)}(\theta_{1},\theta_{2})=\alpha^{(r)}_{e}(e^{r\theta_{2}\ell^{(r)}_{0}}-e^{r\theta_{1}\ell^{(r)}_{0}})\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s-})}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\quad+\alpha^{(r)}_{e}e^{r\theta_{1}\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[(e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s-}}-e^{-\zeta^{(r)}(r\theta_{1})\widehat{R}^{(r)}_{s-}})1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\quad+\alpha^{(r)}_{e}(e^{r\theta_{1}(\ell^{(r)}_{0}+1)}-e^{r\theta_{2}(\ell^{(r)}_{0}+1)})\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{1}(r\theta_{1})\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{0})]$
$\quad+\alpha^{(r)}_{e}e^{r\theta_{2}(\ell^{(r)}_{0}+1)}\mathbb{E}^{(r)}_{s}[(e^{-\eta^{(r)}_{1}(r\theta_{1})\widehat{R}^{(r)}_{e}}-e^{-\eta^{(r)}_{2}(r\theta_{2})\widehat{R}^{(r)}_{e}})1(L^{(r)}(0)=\ell^{(r)}_{0})].$ (4.24)
@@
$\frac{1}{2}\sum_{i=1,2}c_{i}\sigma_{i}^{2}\big{(}(\beta_{i}-\theta_{i})\mu\theta_{i}r^{2}+o((\theta r)^{2})\big{)}\psi^{(r)}_{i}(\theta_{i})$
@@ 4.3 Expansion of the asymptotic BAR
(Step 1)   In this step, we show that the moments of $\widehat{T}^{(r)}_{v}$ for $v=(e,1),(e,2),s$ and $\widehat{R}^{(r)}_{v^{\prime}}$ for $v^{\prime}=e,s$ are well behaved as $r\downarrow 0$. This is also examined for $\widehat{R}^{(r)}_{s-}1(L^{(r)}=\ell^{(r)}_{0})$ and $\widehat{R}^{(r)}_{e}1(L^{(r)}=\ell^{(r)}_{0-})$ under $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{s}$, respectively. These are basis for the expansion of the BAR. We then introduce the set of key quantities, $\Delta^{(r)}_{i}$ for $i=1,2$.
@@
Under the assumption (3.a), we have, for $k=1,2,3$ as $r\downarrow 0$,
$c^{(r)}_{1}k\mathbb{E}[(R^{(r)}_{s})^{k-1}(1-1(R^{(r)}_{s}\geq 1/r))1(1\leq L^{(r)}\leq\ell^{(r)}_{0})]$
$\quad=\alpha^{(r)}_{e}\Big{(}-\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s-})^{k}1(L^{(r)}(0-)=\ell^{(r)}_{0})]+\mathbb{E}[(\widehat{T}^{(r)}_{s})^{k}]\mathbb{P}^{(r)}_{s}[L^{(r)}(0)\leq\ell^{(r)}_{0}]\Big{)},$ (4.30)
$c^{(r)}_{2}k\mathbb{E}[(R^{(r)}_{s})^{k-1}(1-1(R^{(r)}_{s}\geq 1/r))1(L^{(r)}>\ell^{(r)}_{0})]$
$\quad=\alpha^{(r)}_{e}\Big{(}\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s-})^{k}1(L^{(r)}(0-)=\ell^{(r)}_{0})]+\mathbb{E}[(\widehat{T}^{(r)}_{s})^{k}]\mathbb{P}^{(r)}_{s}[L^{(r)}(0)>\ell^{(r)}_{0}]\Big{)},$ (4.31)
$k\mathbb{E}[(R^{(r)}_{e})^{k-1}(1-1(R^{(r)}_{e}\geq 1/r))1(L^{(r)}\leq\ell^{(r)}_{0})]$
$\quad=\alpha^{(r)}_{e}\Big{(}\mathbb{E}_{s}[(\widehat{R}^{(r)}_{e})^{k}1(L^{(r)}(0)=\ell^{(r)}_{0})]+\mathbb{E}[(\widehat{T}^{(r)}_{e,1})^{k}]\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)<\ell^{(r)}_{0}]\Big{)},$ (4.32)
$k\mathbb{E}[(R^{(r)}_{e})^{k-1}(1-1(R^{(r)}_{e}\geq 1/r))1(L^{(r)}>\ell^{(r)}_{0})]$
$\quad=\alpha^{(r)}_{e}\Big{(}-\mathbb{E}_{s}[(\widehat{R}^{(r)}_{e})^{k}1(L^{(r)}(0)=\ell^{(r)}_{0})]+\mathbb{E}[(\widehat{T}^{(r)}_{e,2})^{k}]\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)\geq\ell^{(r)}_{0}]\Big{)}.$ (4.33)
@@
For $k\geq 1$, we first prove (4.5). We apply $f(z,{\boldsymbol{x}})=(x_{2}\wedge 1/r)^{k}1(z\leq\ell^{(r)}_{0})$ to (4.1), then we have (4.5) because
$\frac{d}{dt}f(X^{(r)}(t))=k(R^{(r)}_{s}(t))^{k-1}1(R^{(r)}_{s}<1/r,1\leq L^{(r)}(t)\leq\ell^{(r)}_{0}).$
Similarly, (4.5), (4.5) and (4.5) are obtained from (4.1) by applying $f(z,{\boldsymbol{x}})=(x_{2}\wedge 1/r)^{k}1(z>\ell^{(r)}_{0})$, $f(z,{\boldsymbol{x}})=(x_{1}\wedge 1/r)^{k}1(z\leq\ell^{(r)}_{0})$ and $f(z,{\boldsymbol{x}})=(x_{1}\wedge 1/r)^{k}1(z>\ell^{(r)}_{0})$, respectively. ∎
@@
For each $k=1,2,3$, $\mathbb{E}_{e}[(\widehat{R}^{(r)}_{s-})^{k}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$ and $\mathbb{E}_{s}[(\widehat{R}^{(r)}_{e})^{k}1(L^{(r)}(0)=\ell^{(r)}_{0})]$ are uniformly bounded for $r\in(0,1]$.
@@
$\Delta^{(r)}_{i}$ $=\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s-}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\quad+\lambda^{(r)}_{i}\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]-\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}].$ (4.34)
@@
$\alpha^{(r)}_{e}\Delta^{(r)}_{1}=-rb_{1}\mu\mathbb{P}[L^{(r)}\leq\ell^{(r)}_{0}]+c^{(r)}_{1}\mu^{(r)}\mathbb{P}[L^{(r)}=0]+o(r),$ (4.35)
$\alpha^{(r)}_{e}\Delta^{(r)}_{2}=r\mu b_{2}\mathbb{P}[L^{(r)}>\ell^{(r)}_{0}]+o(r).$ (4.36)

### v8

@@ 4.2 Asymptotic BAR for the pre-limit process $X^{(r)}(\cdot)$
$\varphi^{(r)}_{1}(\theta_{1})=\mathbb{E}[e^{\theta_{1}rL^{(r)}}1(L^{(r)}\leq\ell^{(r)}_{1})],\qquad\varphi^{(r)}_{2}(\theta_{2})=\mathbb{E}[e^{\theta_{2}rL^{(r)}}1(\ell^{(r)}_{1}<L^{(r)})],$
$\varphi^{(r)}_{1|cd}(\theta_{1})=\mathbb{E}[e^{\theta_{1}rL^{(r)}}|L^{(r)}\leq\ell^{(r)}_{1}],\qquad\varphi^{(r)}_{2|cd}(\theta_{2})=\mathbb{E}[e^{\theta_{2}rL^{(r)}}|\ell^{(r)}_{1}<L^{(r)}].$
@@
$w^{(r)}_{1}(z)=1(z\leq\ell^{(r)}_{1}),\qquad w^{(r)}_{2}(z)=1(z>\ell^{(r)}_{1}),\qquad z\in\mathbb{Z}_{+},$
@@
Recall that (3.a)–(3.e) are always assumed in Section 4 if particular assumptions among them are specified. By Lemma 4.2, $g^{(r)}_{i,r\theta_{i}}(R^{(r)})$ is bounded by $e^{|\theta_{i}|(d_{e,i}+d_{s})}$ for sufficiently small $r$ for each $\theta_{1}\in\mathbb{R}$ and $\theta_{2}\leq 0$, so we have the following facts.
@@
because $r\theta_{1}L^{(r)}1(L^{(r)}\leq\ell^{(r)}_{1})\leq\theta_{1}(\ell+o(r))$ for $\theta_{1}>0$. These finiteness guarantees our computations for $\theta_{1}\in\mathbb{R}$ and $\theta_{2}\leq 0$. We also note the following fact.
@@
$\lim_{r\downarrow 0}\left|\psi^{(r)}_{i}(\theta_{i})-\varphi^{(r)}_{i}(\theta_{i})\right|=0,\quad i=1,2.$ (4.18)
@@
$\left|\psi^{(r)}_{i}(\theta_{i})-\varphi^{(r)}_{i}(\theta_{i})\right|$ $=\left|\mathbb{E}[e^{r\theta_{i}L^{(r)}}(g_{1,\theta_{i}}(R^{(r)}_{e})-1)1(L^{(r)}\leq\ell^{(r)}_{1})]\right|$
$\leq e^{|\theta_{1}|r\ell^{(r)}_{1}}\left|\mathbb{E}[g_{i,\theta_{i}}(R^{(r)}_{e})-1]\right|\to 0,\qquad r\downarrow 0.$
@@
Since $g^{(r)}_{i,\theta_{i}}(R^{(r)}(t))$ changes at $t=0$ only when $L^{(r)}(0-)=\ell^{(r)}_{1}$ under $\mathbb{P}^{(r)}_{e}$ or $L^{(r)}(0)=\ell^{(r)}_{1}$ under $\mathbb{P}^{(r)}_{s}$, (4.9) implies that
$\mathbb{E}^{(r)}_{e}\left[\Delta_{e}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)})(0)1(L^{(r)}(0-)\not=\ell^{(r)}_{1})\right]=0,$
$\mathbb{E}^{(r)}_{s}\left[\Delta_{s}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)})(0)1(L^{(r)}(0)\not=\ell^{(r)}_{1})\right]=0.$
Hence, from the definitions of $X^{(r)}_{1}(0)$ and $\Delta_{e}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)})(0)$, we have
$\mathbb{E}^{(r)}_{e}[\Delta_{e}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)})(0)]=\mathbb{E}^{(r)}_{e}[(f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}_{1}(0))-f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0-)))1(L^{(r)}(0-)=\ell^{(r)}_{1})]$
$\quad=\mathbb{E}^{(r)}_{e}\Big{[}e^{r\theta_{2}(\ell^{(r)}_{1}+1)}g^{(r)}_{2,r\theta_{2}}(R^{(r)}_{1}(0))1(L^{(r)}_{1}(0)=\ell^{(r)}_{1}+1)$
$\hskip 64.58313pt-e^{r\theta_{1}\ell^{(r)}_{1}}g^{(r)}_{1,r\theta_{1}}(R^{(r)}(0-)))1(L^{(r)}(0-)=\ell^{(r)}_{1})\Big{]}$
$\quad=e^{r\theta_{2}\ell^{(r)}_{1}}\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s-}}1(L^{(r)}(0-)=\ell^{(r)}_{1})]$
$\hskip 64.58313pt-e^{r\theta_{1}\ell^{(r)}_{1}}\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta_{1})\widehat{R}^{(r)}_{s-}}1(L^{(r)}(0-)=\ell^{(r)}_{1}))],$
$\quad=(e^{r\theta_{2}\ell^{(r)}_{1}}-e^{r\theta_{1}\ell^{(r)}_{1}})\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s-}}1(L^{(r)}(0-)=\ell^{(r)}_{1})]$
$\qquad+e^{r\theta_{1}\ell^{(r)}_{1}}\mathbb{E}^{(r)}_{e}[(e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s-}}-e^{-\zeta^{(r)}(r\theta_{1})\widehat{R}^{(r)}_{s-}})1(L^{(r)}(0-)=\ell^{(r)}_{1}))],$ (4.23)
@@
$\mathbb{E}^{(r)}_{s}[\Delta_{s}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)})(0)]=\mathbb{E}^{(r)}_{s}[(f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0))-f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}_{1}(0)))1(L^{(r)}(0)=\ell^{(r)}_{1})]$
$\quad=\mathbb{E}^{(r)}_{s}\Big{[}e^{r\theta_{1}\ell^{(r)}_{1}}g^{(r)}_{1,r\theta_{1}}(R^{(r)}(0))1(L^{(r)}(0)=\ell^{(r)}_{1})$
$\qquad-e^{r\theta_{2}(\ell^{(r)}_{1}+1)}g^{(r)}_{2,r\theta_{2}}(R^{(r)}_{1}(0)))1(L^{(r)}_{1}(0-)=\ell^{(r)}_{1}+1)\Big{]}$
$\quad=e^{r\theta_{1}(\ell^{(r)}_{1}+1)}\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{1}(r\theta_{1})\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{1})]$
$\qquad-e^{r\theta_{2}(\ell^{(r)}_{1}+1)}\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{2}(r\theta_{2})\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{1})]$
$\quad=(e^{r\theta_{1}(\ell^{(r)}_{1}+1)}-e^{r\theta_{2}(\ell^{(r)}_{1}+1)})\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{1}(r\theta_{1})\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{1})]$
$\qquad+e^{r\theta_{2}(\ell^{(r)}_{1}+1)}\mathbb{E}^{(r)}_{s}[(e^{-\eta^{(r)}_{1}(r\theta_{1})\widehat{R}^{(r)}_{e}}-e^{-\eta^{(r)}_{2}(r\theta_{2})\widehat{R}^{(r)}_{e}})1(L^{(r)}(0)=\ell^{(r)}_{1})].$ (4.24)
Define $D^{(r)}(\theta_{1},\theta_{2})=\alpha^{(r)}_{e}\left(\mathbb{E}^{(r)}_{e}\left[\Delta_{e}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)})(0)\right]+\mathbb{E}^{(r)}_{s}\left[\Delta_{s}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)})(0)\right]\right)$. Then, by (4.2) and (4.2),
$D^{(r)}(\theta_{1},\theta_{2})=\alpha^{(r)}_{e}(e^{r\theta_{2}\ell^{(r)}_{1}}-e^{r\theta_{1}\ell^{(r)}_{1}})\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s-})}1(L^{(r)}(0-)=\ell^{(r)}_{1})]$
$\quad+\alpha^{(r)}_{e}e^{r\theta_{1}\ell^{(r)}_{1}}\mathbb{E}^{(r)}_{e}[(e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s-}}-e^{-\zeta^{(r)}(r\theta_{1})\widehat{R}^{(r)}_{s-}})1(L^{(r)}(0-)=\ell^{(r)}_{1})]$
$\quad+\alpha^{(r)}_{e}(e^{r\theta_{1}(\ell^{(r)}_{1}+1)}-e^{r\theta_{2}(\ell^{(r)}_{1}+1)})\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{1}(r\theta_{1})\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{1})]$
$\quad+\alpha^{(r)}_{e}e^{r\theta_{2}(\ell^{(r)}_{1}+1)}\mathbb{E}^{(r)}_{s}[(e^{-\eta^{(r)}_{1}(r\theta_{1})\widehat{R}^{(r)}_{e}}-e^{-\eta^{(r)}_{2}(r\theta_{2})\widehat{R}^{(r)}_{e}})1(L^{(r)}(0)=\ell^{(r)}_{1})].$ (4.25)
@@
$\frac{1}{2}\sum_{i=1,2}c_{i}\sigma_{i}^{2}\big{(}(\beta_{i}-\theta_{i})\mu\theta_{i}r^{2}+o((\theta_{i}r)^{2})\big{)}\psi^{(r)}_{i}(\theta_{i})$
@@ 4.3 Expansion of the asymptotic BAR
(Step 1)   In this step, we show that the moments of $\widehat{T}^{(r)}_{v}$ for $v=(e,1),(e,2),s$ and $\widehat{R}^{(r)}_{v^{\prime}}$ for $v^{\prime}=e,s$ are well behaved as $r\downarrow 0$. This is also examined for $\widehat{R}^{(r)}_{s-}1(L^{(r)}=\ell^{(r)}_{1})$ and $\widehat{R}^{(r)}_{e}1(L^{(r)}=\ell^{(r)}_{0-})$ under $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{s}$, respectively. These are basis for the expansion of the BAR. We then introduce the set of key quantities, $\Delta^{(r)}_{i}$ for $i=1,2$.
@@
For $k=1,2,3$ and $r\in(0,1]$,
$c^{(r)}_{1}k\mathbb{E}[(R^{(r)}_{s})^{k-1}(1-1(R^{(r)}_{s}\geq 1/r))1(1\leq L^{(r)}\leq\ell^{(r)}_{1})]$
$\quad=\alpha^{(r)}_{e}\Big{(}-\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s-})^{k}1(L^{(r)}(0-)=\ell^{(r)}_{1})]+\mathbb{E}[(\widehat{T}^{(r)}_{s})^{k}]\mathbb{P}^{(r)}_{s}[L^{(r)}(0)\leq\ell^{(r)}_{1}]\Big{)},$ (4.31)
$c^{(r)}_{2}k\mathbb{E}[(R^{(r)}_{s})^{k-1}(1-1(R^{(r)}_{s}\geq 1/r))1(L^{(r)}>\ell^{(r)}_{1})]$
$\quad=\alpha^{(r)}_{e}\Big{(}\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s-})^{k}1(L^{(r)}(0-)=\ell^{(r)}_{1})]+\mathbb{E}[(\widehat{T}^{(r)}_{s})^{k}]\mathbb{P}^{(r)}_{s}[L^{(r)}(0)>\ell^{(r)}_{1}]\Big{)},$ (4.32)
$k\mathbb{E}[(R^{(r)}_{e})^{k-1}(1-1(R^{(r)}_{e}\geq 1/r))1(L^{(r)}\leq\ell^{(r)}_{1})]$
$\quad=\alpha^{(r)}_{e}\Big{(}\mathbb{E}_{s}[(\widehat{R}^{(r)}_{e})^{k}1(L^{(r)}(0)=\ell^{(r)}_{1})]+\mathbb{E}[(\widehat{T}^{(r)}_{e,1})^{k}]\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)<\ell^{(r)}_{1}]\Big{)},$ (4.33)
$k\mathbb{E}[(R^{(r)}_{e})^{k-1}(1-1(R^{(r)}_{e}\geq 1/r))1(L^{(r)}>\ell^{(r)}_{1})]$
$\quad=\alpha^{(r)}_{e}\Big{(}-\mathbb{E}_{s}[(\widehat{R}^{(r)}_{e})^{k}1(L^{(r)}(0)=\ell^{(r)}_{1})]+\mathbb{E}[(\widehat{T}^{(r)}_{e,2})^{k}]\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)\geq\ell^{(r)}_{1}]\Big{)}.$ (4.34)
@@
For $k\geq 1$, we first prove (4.5). We apply $f(z,{\boldsymbol{x}})=(x_{2}\wedge 1/r)^{k}1(z\leq\ell^{(r)}_{1})$ to (4.1), then we have (4.5) because
$\frac{d}{dt}f(X^{(r)}(t))=k(R^{(r)}_{s}(t))^{k-1}1(R^{(r)}_{s}<1/r,1\leq L^{(r)}(t)\leq\ell^{(r)}_{1}).$
Similarly, (4.5), (4.5) and (4.5) are obtained from (4.1) by applying $f(z,{\boldsymbol{x}})=(x_{2}\wedge 1/r)^{k}1(z>\ell^{(r)}_{1})$, $f(z,{\boldsymbol{x}})=(x_{1}\wedge 1/r)^{k}1(z\leq\ell^{(r)}_{1})$ and $f(z,{\boldsymbol{x}})=(x_{1}\wedge 1/r)^{k}1(z>\ell^{(r)}_{1})$, respectively. ∎
@@
For each $k=1,2,3$, $\mathbb{E}_{e}[(\widehat{R}^{(r)}_{s-})^{k}1(L^{(r)}(0-)=\ell^{(r)}_{1})]$ and $\mathbb{E}_{s}[(\widehat{R}^{(r)}_{e})^{k}1(L^{(r)}(0)=\ell^{(r)}_{1})]$ are uniformly bounded for $r\in(0,1]$.
@@
$\Delta^{(r)}_{i}$ $=\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s-}1(L^{(r)}(0-)=\ell^{(r)}_{1})]$
$\quad+\lambda^{(r)}_{i}\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{1})]-\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{1}].$ (4.35)
@@
$\alpha^{(r)}_{e}\Delta^{(r)}_{1}=-rb_{1}\mu\mathbb{P}[L^{(r)}\leq\ell^{(r)}_{1}]+c^{(r)}_{1}\mu^{(r)}\mathbb{P}[L^{(r)}=0]+o(r),$ (4.36)
$\alpha^{(r)}_{e}\Delta^{(r)}_{2}=r\mu b_{2}\mathbb{P}[L^{(r)}>\ell^{(r)}_{1}]+o(r).$ (4.37)